If the genotypes A1A1, A1A2, and A2A2 have frequencies 0.5, 0.25, and 0.25 respectively, what are the frequencies of the A1 and A2 alleles? b) After a single generation of random mating, what are the expected frequency of A1A1, A1A2 and A2A2 in the offspring? c) If A2 is a recessive allele causing freckles, how many freckled offspring do you expect if there are 441 offspring in total?
Answer #1

The survival to sexual maturity rates of genotype A1A1 is 90%, A1A2 is 85% and A2A2 is 75%. The fecundity of genotype A1A1 is 50, A1A2 is 55 and A2A2 is 70.

So, the absolute fitness i.e W of these genotypes can be determined by multiplying their survival rates by their fecundities.

  • Absolute fitness of genotype A1A1 = 90/100 × 50 = 45
  • Absolute fitness of genotype A1A2 = 85/100 × 55 = 46.75
  • Absolute fitness of genotype A2A2 = 75/100 × 70 = 52.5

Therefore, the absolute fitness for genotypes A1A1, A1A2, A2A2 are 45, 46.75 and 52.5 respectively.

(b) By taking fitness reference of genotype A1A1 i.e 45, the relative fitness of the genotypes can be determined by dividing the absolute fitness of the genotypes by fitness reference.

  • Relative fitness of genotype A1A1 = 45/45 = 1
  • Relative fitness of genotype A1A2 = 46.75/45 = 1.03
  • Relative fitness of genotype A2A2 = 52.5/45 = 1.16

Therefore, the relative fitness for genotypes A1A1, A1A2, A2A2 are 1, 1.03, 1.16 respectively.

(c) If the frequency of the A2 allele is p = 0.5. This frequency will remain same after one generation without the presence of evolutionary factors according to Hardy Weinberg principle.